# Stability of spherical thin-shell wormholes in scalar-tensor theories

### v2

@@ 1 Introduction
In the context of the different theories of gravity, there are solutions of the field equations that represent compact objects, such as black holes and wormholes. The existence of wormholes is still hypothetical; they were first introduced by Einstein and Rosen [23] and studied more systematically since the work by Morris and Thorne [24]. These geometries represent bridges that connect two different regions of the Universe or two different universes across a throat [25], where the flare-out condition is fulfilled. A common issue encountered in traversable wormholes is the presence of the so-called exotic matter, i.e. matter that does not satisfy the energy conditions [25], which is not a physically desirable feature. In GR, the exotic matter is always required near the throat [26, 27, 28]; however, its amount can be reduced [29], but at the cost of increasing tensions in it [30]. Many papers concerning wormholes with scalar fields have appeared in the literature, here we can only mention a few [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. On the other hand, a new spacetime can be obtained by the cut and paste of two geometries, which are joined on their boundaries by applying the suitable junction conditions, firstly found in GR [43, 44]. This method is used to construct the denominated thin-shell wormholes [25], that allows to localize and minimize the amount of exotic matter. With this procedure, the new manifold typically has an infinitesimally thin layer of matter at the joining hypersurface, where the throat is placed. The construction, the stability, and other related topics on thin-shell wormholes have been studied within GR [45, 46, 47, 48, 49, 50, 51], in dilaton gravity [52], in Einstein-Gauss-Bonnet [53, 54], in Brans-Dicke [55], in $f(R)$ theory [56, 57, 58, 59, 60, 61, 62], and also in Palatini $f(R)$ gravity [63], among others.
@@
In this paper, we construct thin-shell wormholes with spherical symmetry in Einstein theory non-minimally coupled to a scalar field. In order to have a proper matching at the shell, we apply the junction conditions corresponding to this theory [64]. For a family of wormholes symmetric across the throat, which is located at the shell, we study their dynamical stability and the presence of exotic matter. In particular, for our construction we adopt a black hole solution within Einstein-Maxwell gravity coupled to a conformally invariant scalar field, having a constant scalar field [17]. The structure of this paper is as follows. In Sect. 2, we present the general construction of the wormholes in scalar-tensor gravity. In Sect. 3, we develop the formalism for the stability of wormholes symmetric across the throat. In Sect. 4, we consider a particular example of an spacetime with a radial electric field, in which we study the stability and the energy conditions. Finally, we summarize our results in Sect. 5. We use units such that $c=1$.
@@ 3 Stability of wormholes symmetric across the throat
We look for an expression of $\dot{a}^{2}$ in the form of Eq. (25). With the assumption $f^{\prime}(\phi)\neq 0$, the help of the definitions
@@ 4 Wormholes with a radial electric field
We present our results graphically. In Fig. 1, we study the existence and stability of thin-shell wormholes by displaying $a_{0}/m$ against $Q/m$ for different values of the parameter $s/m^{2}$. The black curve corresponds to the solutions that satisfy (33), where the stable configurations are drawn with a solid line and the unstable ones with a dashed line. The dark gray areas show the regions without physical meaning which are removed, while the light gray ones are those where the scalar field is not real. In more detail, for negative $s/m^{2}$, represented by the plot with $s/m^{2}=-1$, we find two solutions for $Q>Q_{c}$, the larger one is stable, while the other is unstable, but in both the scalar field is not real so they have to be discarded. For $0<s/m^{2}<1$, represented by $s/m^{2}=0.5$, two solutions are found when $Q>Q_{c}$, a stable, larger one and an unstable, smaller one. The region without physical meaning shrinks to finally disappear as we increase the value of $s/m^{2}$. For $s/m^{2}=1$, there are two solutions for small values of $Q/m^{2}$; the larger one is stable while the other is unstable. In this case, there are no horizons in the original metric for $Q\neq 0$, and for the case of $Q=0$, there is a horizon where the solutions fall on its border or into it. For $1<s/m^{2}\leq 9/8$, represented by $s/m^{2}=1.1$ the behavior is similar to the case with $s/m^{2}=1$. Finally, if $s/m^{2}\geq 9/8$, one can see that Eq. (33) has no real solutions, so our construction is not possible.

### v3

@@ 1 Introduction
In the context of the different theories of gravity, there are solutions of the field equations that represent compact objects, such as black holes and wormholes. The existence of wormholes is still hypothetical; they were first introduced by Einstein and Rosen [23] and studied more systematically since the work by Morris and Thorne [24]. These geometries represent bridges that connect two different regions of the Universe or two different universes across a throat [25], where the flare-out condition is fulfilled. A common issue encountered in traversable wormholes is the presence of the so-called exotic matter, i.e. matter that does not satisfy the energy conditions [25], which is not a physically desirable feature. In GR, the exotic matter is always required near the throat [26, 27, 28]; however, its amount can be reduced [29], but at the cost of increasing tensions in it [30]. Many papers concerning wormholes with scalar fields have appeared in the literature, here we can only mention a few [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. On the other hand, a new spacetime can be obtained by the cut and paste of two geometries, which are joined on their boundaries by applying the suitable junction conditions, firstly found in GR [43, 44]. This method is used to construct the denominated thin-shell wormholes [25], in which the new manifold has an infinitesimally thin layer of matter at the joining hypersurface, where the throat is placed. This procedure allows to localize and minimize the exotic matter at the throat. Although matter not satisfying the energy conditions appears in many physical contexts, in the case of wormholes is advantageous to reduce the amount required at the throat. Another interesting aspect of thin-shell wormholes is the possibility of performing analytical studies of stability in the case of highly symmetric static configurations, under perturbations preserving the symmetry. The construction, the stability, and other related topics on thin-shell wormholes have been considered within GR [45, 46, 47, 48, 49, 50, 51], in dilaton gravity [52], in Einstein-Gauss-Bonnet [53, 54], in Brans-Dicke [55], in $f(R)$ theory [56, 57, 58, 59, 60, 61, 62], and also in Palatini $f(R)$ gravity [63], among others.
@@
In this paper, we construct thin-shell wormholes with spherical symmetry in Einstein theory non-minimally coupled to a scalar field. In order to have a proper matching at the shell, we apply the junction conditions corresponding to this theory [64, 65]. For a family of wormholes symmetric across the throat, which is located at the shell, we study their dynamical stability and the presence of exotic matter. In particular, for our construction we adopt a black hole solution within Einstein-Maxwell gravity coupled to a conformally invariant scalar field, having a constant scalar field [17]. The structure of this paper is as follows. In Sect. 2, we present the general construction of the wormholes in scalar-tensor gravity. In Sect. 3, we develop the formalism for the stability of wormholes symmetric across the throat. In Sect. 4, we consider a particular example of a spacetime with a radial electric field, in which we study the stability and the energy conditions. Finally, we summarize our results in Sect. 5. We use units such that $c=1$.
@@ 3 Stability of wormholes symmetric across the throat
We look for an expression of $\dot{a}^{2}$ in the form of Eq. (25). As we are considering a non-minimally coupled scalar field, we assume that $f^{\prime}(\phi)\neq 0$. With the help of the definitions
@@ 4 Wormholes with a radial electric field
We present our results graphically. In Fig. 1, we study the existence and stability of thin-shell wormholes by displaying $a_{0}/m$ against $Q/m$ for different values of the parameter $s/m^{2}$. The black curve corresponds to the solutions that satisfy (33), where the stable configurations are drawn with a solid line and the unstable ones with a dashed line. The dark gray areas show the regions without physical meaning which are removed, while the light gray ones are those where the scalar field is not real. In more detail, for negative $s/m^{2}$, represented by the plot with $s/m^{2}=-1$, we find two solutions for $Q>Q_{c}$, the larger one is stable, while the other is unstable, but in both the scalar field is not real so they have to be discarded. For $0<s/m^{2}<1$, represented by $s/m^{2}=0.5$, two solutions are found when $Q>Q_{c}$, a stable, larger one and an unstable, smaller one. The region without physical meaning shrinks to finally disappear as we increase the value of $s/m^{2}$. For $s/m^{2}=1$, there are two solutions for small values of $Q/m^{2}$; the larger one is stable while the other is unstable. In this case, there are no horizons in the original metric for $Q\neq 0$, and for the case of $Q=0$, there is a horizon where the solutions fall on its border or into it. For $1<s/m^{2}<9/8$, represented by $s/m^{2}=1.1$ the behavior is similar to the case with $s/m^{2}=1$. Finally, if $s/m^{2}\geq 9/8$, one can see that Eq. (33) has no real solutions, so our construction is not possible.